# LightStereo: Channel Boost Is All Your Need for Efficient 2D Cost Aggregation

### v1

@@ 3.3 Network Architecture of LightStereo
Our proposed LightStereo consists of five components: feature extraction, cost computation, cost aggregation, disparity prediction, and loss. In the following, we provide introductions to each module.
@@
Multi-scale Feature Extraction. For a stereo image pair input with dimensions H × W × 3, our methodology harnesses a MobileNetV2 bangunharcana2021coex ; xu2023iterative model, previously trained on the ImageNet deng2009imagenet dataset, serving as the core architectural foundation. We enable the extraction of feature maps across four distinct scales, effectively reducing the resolution to 1/4, 1/8, 1/16, and 1/32 of the initial size, respectively. Subsequently, upsampling blocks with skip connections are utilized to restore these feature maps to 1/4 scale, which is used to construct the cost volume.
@@
Cost Volume. Utilizing the left features $\mathbf{f}_{l,4}$ and right features $\mathbf{f}_{r,4}$ extracted from images $\mathbf{I}_{l}$ and $\mathbf{I}_{r}$, we construct a correlation volume. The correlation cost volume construction involves comparing features from $\mathbf{f}_{l,4}$ and $\mathbf{f}_{r,4}$ across different disparity levels. For each disparity $d$ within the range of 0 to $D-1$, the similarity between $\mathbf{f}_{l,4}$ and $\mathbf{f}_{r,4}$ shifted by $d$ pixels is computed. Mathematically, this can be expressed as:
@@
Cost Aggregation. We use inverted residual blocks to aggregate the cost volume at 1/4, 1/8, and 1/16 resolutions, and at each resolution, we apply multi-scale convolutional attention using the left image, as described in Section 3.1 and Section 3.2. We develop three variants of LightStereo based on block and expansion variations: LightStereo-S for small-scale applications, LightStereo-M for medium-scale tasks, and LightStereo-L. In LightStereo-S, the inverted residual blocks are configured as (1, 2, 4) with an expansion factor of 4. For LightStereo-M, the blocks are set as (4, 8, 14) with an expansion factor of 4. Finally, LightStereo-L utilizes blocks (8, 16, 32) with an expansion factor of 8. LightStereo-L denotes that the backbone of LightStereo-L has been replaced with EfficientnetV2 tan2021efficientnetv2 .
Disparity Regression. We utilize disparity regression gcnet ; psmnet2018 to estimate the disparity map. This method predicts disparities by leveraging the probability distribution of each disparity $d$, derived from the predicted cost $c_{d}$ through the softmax operation $\sigma(\cdot)$. The predicted disparity $\hat{d}$ is then determined by summing each disparity $d$ weighted by its probability, as depicted by the equation:
@@
Loss We employ the smooth $L_{1}$ loss to train the proposed LightStereo. The loss is defined as:
@@
4 Experiment
4.1 Datasets and Evaluation Metrics
SceneFlowsceneflow is a synthetic stereo collection that provides 35,454 and 4,370 image pairs for training and testing, respectively. The dataset has a resolution of 960×540 and uses dense disparity maps as ground truth. The data is split into two categories: cleanpass and finalpass. Cleanpass refers to the synthetic images that are generated by clean renderings without post-processing, while finalpass images are produced with photorealistic settings such as motion blur, defocus blur, and noise. In evaluations, we utilize the widely-used metrics the end point error (EPE) as the evaluation metrics.
KITTI. KITTI 2012kitti2012 and KITTI 2015kitti2015 are datasets that captured from real-world scenes. KITTI 2012 contains 194 and 195 image pairs for training and testing, respectively. KITTI 2015 provides two sets with 200 image pairs for training and testing. All KITTI datasets provide sparse ground-truth disparities guided by the LiDAR system. For evaluations, we calculate EPE and the percentage of pixels with EPE larger than 3 pixels in all (D1-all) regions. All two KITTI datasets are also used for cross-domain generalization performance evaluation, with EPE and a >3px metric (i.e., the percentage of points with absolute error larger than 3 pixels) reported.
4.2 Implementation Details
We have implemented our methods using PyTorch and conducted experiments on 8 NVIDIA RTX 3090 GPUs. For the SceneFLow sceneflow datasets, the batch sizes for LightStereo-S, LightStereo-M, LightStereo-L, and LightStereo-L are set at 24, 12, 8, and 6, respectively. We utilize the AdamW optimizer coupled with OneCycleLR scheduling, where the maximum learning rate was set to 0.0001 multiplied by the batch size. In the ablation study, these LightStereo models were trained over 50 epochs. For the final model evaluation, LightStereo underwent training for 90 epochs. The random crop was used for data augmentation. For the KITTI dataset, we fine-tuned the pre-trained models on the SceneFlow dataset sceneflow for 500 epochs using a mixed training set comprising KITTI 2012 kitti2012 and KITTI2015 kitti2015 training datasets. We employed a batch size of 2 and utilized the AdamW optimizer. OneCycleLR scheduling was used with a max learning rate of 0.0002. Data augmentation techniques including color jitter, random erase, random scale, and random crop were employed.
Figure 4: Qualitative results on the test set of Sceneflow sceneflow . The first column presents the left image on the first row and the right image on the second row. Columns 2 to 5 display the predicted disparity maps on the first row and the corresponding error maps on the second row.
Table 1: Comparison with the state-of-the-art on SceneFlowsceneflow . The runtime is tested on our RTX 3090 GPU.
Method FLOPs(G) Params(M) EPE (px) Runtime (ms)
DeepPruner-Fastdeeppruner2019 219.12 7.47 1.25 39
StereoNetstereonet2018 85.93 0.40 1.10 20
2D-MobileStereoNetshamsafar2022mobilestereonet 128.84 2.23 1.11 73
FADNet++wang2021fadnet++ 148.21 124.26 0.85 21
CoExbangunharcana2021coex 53.39 2.72 0.67 36
Fast-ACVNet fast-acv 79.34 3.08 0.64 22
Fast-ACVNet+ fast-acv 93.08 3.20 0.59 27
AANetxu2020aanet 152.86 2.97 0.87 93
IINetli2024iinet 90.16 19.54 0.54 26
@@ 4.2 Implementation Details
LightStereo-L(Ours) 159.26 45.63 0.51 54
Table 2: Comparison with the state-of-the-art methods on KITTI benchmarks. The methods are categorized based on whether its runtime exceeds 100ms. Numbers in bold represent the highest values, while numbers underlined indicate the second-highest values. The runtime of the model with is tested on our RTX 3090 GPU.
Target Method KITTI 2012kitti2012 KITTI 2015kitti2015
3-noc 3-all 4-noc 4-all EPE noc EPE all D1-bg D1-fg D1-all Runtime (ms)
Accuracy GANetganet2019 1.19 1.60 0.91 1.23 0.4 0.5 1.48 3.46 1.81 1800
LaC+GANetLaC+GANet 1.05 1.42 0.80 1.09 0.4 0.5 1.44 2.83 1.67 1800
CFNetcfnet2021 1.23 1.58 0.92 1.18 0.4 0.5 1.54 3.56 1.88 180
SegStereosegstereo2018 1.68 2.03 1.25 1.52 0.5 0.6 1.88 4.07 2.25 600
SSPCVNet SSPCVNet 1.47 1.90 1.08 1.41 0.5 0.6 1.75 3.89 2.11 900
EdgeStereo-V2song2020edgestereo 1.46 1.83 1.07 1.34 0.4 0.5 1.84 3.30 2.08 320
CSPNCSPN 1.19 1.53 0.93 1.19 - - 1.51 2.88 1.74 1000
LEAStereoleastereo 1.13 1.45 0.83 1.08 0.5 0.5 1.40 2.91 1.65 300
CREStereoCrestereo 1.14 1.46 0.90 1.14 0.4 0.5 1.45 2.86 1.69 410
ACVNet acvnet 1.13 1.47 0.86 1.12 0.4 0.5 1.37 3.07 1.65 200
Speed DispNetCsceneflow 4.11 4.65 2.77 3.20 0.9 1.0 4.32 4.41 4.34 60
StereoNet stereonet2018 - - - - 0.8 0.9 4.30 7.45 4.83 22
DeepPrunerFastdeeppruner2019 - - - - - - 2.32 3.91 2.59 50
AANetxu2020aanet 1.91 2.42 1.46 1.87 0.5 0.6 1.99 5.39 2.55 62
DecNetDecNet - - - - - - 2.07 3.87 2.37 50
HITNethitnet 1.41 1.89 1.14 1.53 0.4 0.5 1.74 3.20 1.98 54
CoExbangunharcana2021coex 1.55 1.93 1.15 1.42 0.5 0.5 1.79 3.82 2.13 33
Fast-ACVNetfast-acv 1.68 2.13 1.23 1.56 0.5 0.6 1.82 3.93 2.17 39
Fast-ACVNet+fast-acv 1.45 1.85 1.06 1.36 0.5 0.5 1.70 3.53 2.01 45
LightStereo-S (Ours) 1.88 2.34 1.30 1.65 0.6 0.6 2.00 3.80 2.30 17
LightStereo-M (Ours) 1.56 1.91 1.10 1.36 0.5 0.5 1.81 3.22 2.04 23
LightStereo-L (Ours) 1.55 1.87 1.10 1.33 0.5 0.5 1.78 2.64 1.93 34
LightStereo-L(Ours) 1.34 1.62 0.96 1.17 0.5 0.5 1.60 2.92 1.82 49
4.3 Comparisons with State-of-the-art Methods
As shown in Table 1, we compare our proposed LightStereo methods with several state-of-the-art stereo matching approaches on the SceneFlow sceneflow dataset. In terms of runtime, LightStereo-S achieves a runtime of only 17ms. This is substantially lower than other methods. Regarding model complexity, LightStereo-S strikes a favorable balance with only 22.71Gflops, comparable to StereoNet stereonet2018 (85.93Gflops). This reflects our commitment to maintaining a lightweight model while ensuring competitive performance. In terms of accuracy, LightStereo-L achieves an EPE of 0.51, which although slightly higher than some methods like Fast-ACVNet+ fast-acv (0.59) and IINet li2024iinet (0.54), remains competitive while offering significantly lower computational costs. Overall, our LightStereo framework, particularly the LightStereo-S configuration, presents a compelling solution for real-time stereo matching, offering a favorable trade-off between computational efficiency and depth estimation accuracy. This makes it highly suitable for applications such as autonomous navigation and augmented reality, where rapid and reliable depth perception is essential. Figure 4 presents the visualization results of the four proposed models on the SceneFlow sceneflow dataset. Additionally, we evaluated the results of our proposed models on the KITTI 2012 and 2015 benchmarks. As illustrated in Table 2, the runtime of LightStereo-S exceeds all methods. Notably, LightStereo-L surpasses all other lightweight state-of-the-art stereo matching networks across every evaluation metric on the KITTI12 dataset. Additionally, LightStereo-L achieved top performance on the KITTI15 dataset in terms of D1-bg and D1-all metrics among lightweight stereo matching methods.
4.4 Ablation Study
@@
Table 3: Ablation study of backbone on SceneFLow sceneflow .
@@
MobilenetV2 sandler2018mobilenetv2 CNN 0.7144 35.82 7.54 22.93
MobilenetV3 mobilenetv3 CNN 0.7292 35.72 9.16 25.02
StarNet StarNet CNN 0.7247 40.21 8.98 26.63
EfficientnetV2 tan2021efficientnetv2 CNN 0.6130 103.14 28.87 46.83
RepVIT wang2023repvit Transformer 0.6823 50.45 9.56 28.65
Backbone. This paper aims to design a lightweight stereo-matching model. To achieve this, we explored the use of classic lightweight models as backbones. As illustrated in Table 3, MobilenetV2 sandler2018mobilenetv2 demonstrated a balanced performance with an EPE of 0.7144, FLOPs of 35.82G, 7.54M parameters, and an inference time of 22.93ms. Although MobilenetV3 mobilenetv3 had slightly lower FLOPs at 35.72G, it showed a higher EPE of 0.7292 and required more parameters (9.16M) with a longer inference time (25.02ms). StarNet StarNet exhibited an EPE of 0.7247 with higher FLOPs (40.21G) and inference time (26.63ms). EfficientnetV2 tan2021efficientnetv2 achieved the lowest EPE of 0.6130, but at the cost of significantly higher computational resources (103.14G FLOPs) and parameters (28.87M), with an inference time of 46.83ms. RepVIT wang2023repvit , a transformer-based model, showed an EPE of 0.6876 but required 50.45GFLOPs and 9.56M parameters with an inference time of 28.65ms. Therefore, MobilenetV2 is chosen for its overall efficiency and balance between accuracy and computational cost.
Figure 5: Qualitative results on the test set of Sceneflow sceneflow . The first column presents the left image on the first row and the right image on the second row. Columns 2 to 5 display the predicted disparity maps on the first row and the corresponding error maps on the second row.
Table 4: Ablation study on the design of the block on SceneFLow sceneflow . Regular represents regular convolution. V1 Block represents the depthwise separable convolution mijwil2023mobilenetv1 . V2 Block represents an inverted residual block sandler2018mobilenetv2 . DW refers to depthwise separable convolution. ViT Block refers to the block used in EfficientViT liu2023efficientvit . The underlined backbone is the one we ultimately used.
Conv. Type Kernel Block EPE (px) Flops(G) Param(M) time(ms)
Regular $3\times 3$ (4 8 16) 0.7652 36.27 8.04 22.75
Regular $5\times 5$ (4 8 16) 0.7979 71.16 18.62 17.61
Regular $7\times 7$ (4 8 16) 0.8190 123.51 34.49 20.06
Regular $11\times 1$ 11 (4 8 16) 0.8672 280.53 82.10 34.22
V1 Block DW $3\times 3$ (30 60 120) 0.7801 34.86 7.57 54.21
V2 Block DW $3\times 3$ (4 8 16) 0.7144 35.82 7.54 22.93
ViT Block - (3 6 9) 0.7149 34.48 6.53 51.14
Conv. Type. As illustrated in Table 4, our experiments highlight the critical importance of disparity dimensions in the cost aggregation process of stereo matching, rather than spatial expansions in the height and width dimensions. Regular convolutions with different kernel sizes showed that increasing kernel size resulted in higher EPE and significantly increased computational cost. Specifically, expanding the receptive field in the height and width dimensions did not significantly contribute to performance improvement. This finding indicates that merely increasing the spatial extent of convolutions does not effectively enhance the accuracy of stereo-matching. Instead, the emphasis should be placed on optimizing the disparity dimensions to improve the aggregation of cost volumes, as this approach is more effective in capturing relevant features and enhancing overall performance. The V1 block with a $3\times 3$ kernel showed an EPE of 0.7801, indicating moderate performance with reduced FLOPs (34.86G), but it had a higher inference time (54.21ms). In contrast, the V2 block provided the best balance, achieving the lowest EPE of 0.7144, and FLOPs of 35.82G, with an inference time of 22.93ms. This superior performance is attributed to the structure of the V2 block, which incorporates expansion convolutions in the disparity dimension. By focusing on enhancing the disparity dimensions rather than merely expanding the spatial dimensions, the V2 block effectively aggregates cost volumes, leading to improved accuracy and efficiency in stereo-matching tasks. The ViT block showed a similar EPE to the V2 block but had a lower parameter count (6.53M) and higher inference time (51.14ms). Ultimately, the V2 block was chosen for its optimal balance between accuracy and computational efficiency. Figure 5 shows a visual comparison with different conv. type. It can be observed that for the occluded area in the lower left corner of the left image, the cost aggregation based on the V2 block achieves better results than the other 3 conv. type.
Block Structure Analysis. As illustrated in Table 5, we explore the impact of different block structures while maintaining a constant expansion factor of 4. Configurations ’e’ to ’h’ explore blocks (1, 2, 4), (2, 4, 8), (4, 8, 16), and (8, 16, 32) respectively. The results show that larger blocks generally lead to better performance. For instance, configuration ’e’ with the smallest block (1, 2, 4) has the highest EPE of 0.8317, while configuration ’h’ with the largest block (8, 16, 32) shows an improvement in EPE, but detailed metrics are not provided. There is a trade-off between accuracy and computational cost, as larger blocks increase FLOPs and parameters.
Expansion Factor Analysis. As illustrated in Table 5, we further analyze the effect of varying expansion factors while keeping the block structure constant. As the expansion factor increases, a consistent decrease in EPE is observed, indicating improved accuracy. Specifically, configuration ’a’ with an expansion factor of 2 has an EPE of 0.7557, while configuration ’d’ with an expansion factor of 16 achieves the lowest EPE of 0.6650. The analysis further proves that the critical importance of disparity dimensions cannot be overstated in the cost aggregation process of stereo matching. The disparity dimension is pivotal in effectively aggregating cost volumes, leading to more accurate and reliable depth estimations. However, this improvement in accuracy comes at the cost of significantly increased computational complexity and parameters, with flops increasing from 26.23G to 93.34G and the inference time increases from 22.44ms to 36.67ms.
Table 5: Ablation study on the design of the block on SceneFLow sceneflow . Expansion refers to the expansion factor. SE represents Squeeze-and-Excitation hu2018squeeze . MSCA represents Multi-Scale Convolutional Attention.
Block Expansion SE MSCA EPE (px) Flops(G) Param(M) time(ms)
@@
MSCA Module Analysis. As illustrated in Table 5, the baseline configuration (i) with blocks (4, 8, 16) achieved an EPE of 0.7144. Incorporating MSCA (configuration k) reduced the EPE further to 0.6809 with minimal changes in FLOPs and parameters. This suggests that MSCA provides the most significant improvement in accuracy with minimal impact on computational efficiency. We also explored the use of the SE module hu2018squeeze . Configurations ’i’ to ’j’ show that although the SE module improves accuracy, reducing EPE from 0.7144 to 0.7036. However, this improvement comes with an increase in parameters from 7.54M to 12.76M and a slight increase in FLOPs and inference time.
Table 6: Runtime breakdown.
@@
LightStereo-L 27.17 1.98 23.64 1.49 54.28
4.5 Runtime Analysis

### v3

@@ III-C LightStereo – Network Architecture
We now describe the overall LightStereo architecture, which consists of four components: feature extraction, cost computation, cost aggregation, and disparity prediction.
@@
Multi-scale Feature Extraction. For a stereo image pair with dimensions $H\times W\times 3$, our network harnesses a MobileNetV2 [39, 30] model, previously trained on the ImageNet [41] dataset. We enable the extraction of feature maps across four distinct scales, effectively reducing the resolution to $\frac{1}{4}$, $\frac{1}{8}$, and $\frac{1}{16}$ and $\frac{1}{32}$ of the initial size, respectively. Subsequently, upsampling blocks with skip connections are utilized to restore these feature maps to $\frac{1}{4}$ scale, which is used to construct the cost volume.
@@
Cost Volume. We construct a correlation volume from the previous left $\mathbf{f}_{l,4}$ and right features $\mathbf{f}_{r,4}$. For each disparity $d$ within the range of 0 to $D-1$, the similarity between $\mathbf{f}_{l,4}$ and $\mathbf{f}_{r,4}$ shifted by $d$ pixels is computed as:
@@
Cost Aggregation. We use inverted residual blocks to aggregate the cost volume at $\frac{1}{4}$, $\frac{1}{8}$, and $\frac{1}{16}$ resolutions, and at each resolution, we apply multi-scale convolutional attention using the left image, as described in Section III-A and Section III-B. We develop four variants of LightStereo based on block and expansion variations: LightStereo-S for small-scale applications, LightStereo-M for medium-scale tasks, LightStereo-L for large-scale tasks, and LightStereo-H for huge-scale tasks. In LightStereo-S, the inverted residual blocks are configured as (1, 2, 4) with an expansion factor of 4. For LightStereo-M, the blocks are set as (4, 8, 14) with an expansion factor of 4. Finally, LightStereo-L uses blocks (8, 16, 32) with an expansion factor of 8. LightStereo-H denotes a LightStereo-L variant with EfficientnetV2 [42] backbone.
Disparity Regression. We use soft-argmax [1, 2] to predict the final disparity map $\hat{d}$, determined by summing each disparity $d$ weighted by its probability $\sigma(c_{d})$, with $c_{d}$ being the predicted cost and $\sigma(\cdot)$ a softmax layer:
@@
Training Loss. We train LightStereo with smooth $L_{1}$ loss
@@
TABLE I: Results on SceneFlow [25]. Comparison with state-of-the-art models pursuing speed over accuracy.
Method FLOPs(G) Params(M) EPE (px) Time(ms)
DeepPruner-Fast[34] 219.12 7.47 1.25 39
StereoNet[43] 85.93 0.40 1.10 20
2D-MobileStereoNet[17] 128.84 2.23 1.11 73
FADNet++[16] 148.21 124.26 0.85 21
CoEx[39] 53.39 2.72 0.67 36
Fast-ACVNet [44] 79.34 3.08 0.64 22
Fast-ACVNet+ [44] 93.08 3.20 0.59 27
AANet[18] 152.86 2.97 0.87 93
HITNet [45] 50.23 0.42 0.55 36
IINet[46] 90.16 19.54 0.54 26
@@
LightStereo-H(Ours) 159.26 45.63 0.51 54
IV Experiments
IV-A Datasets and Evaluation Metrics
SceneFlow[25] is a synthetic stereo collection that provides 35 454 and 4 370 image pairs for training and testing, respectively. The dataset has a resolution of 960$\times$540 and provides dense disparity maps as ground truth. Here, we use the End-Point-Error (EPE) as the evaluation metric.
KITTI. KITTI 2012 [26] and 2015 [27] are real-world datasets, counting 194/195 and 200/200 image pairs for training and testing, respectively, for which sparse ground-truth disparities are collected with a LiDAR system. We both report the results obtained from the online benchmarks, for which we report the official metrics, as well as assess cross-domain generalization performance, measuring the percentage of pixels with error >3px in this latter case.
Middlebury 2014. [47] It collects high-resolution indoor scenes, providing 15 image pairs for training and 15 for testing, respectively. We involved this dataset for cross-domain generalization experiments, using half-resolution images and measuring the percentage of pixels with error >2px.
IV-B Implementation Details
LightStereo is implemented in PyTorch and trained on 8 NVIDIA RTX 3090 GPUs. For the SceneFlow [25] dataset, the batch sizes for LightStereo-S, LightStereo-M, LightStereo-L, and LightStereo-H are set at 24, 12, 8, and 6, respectively.We use the AdamW optimizer coupled with OneCycleLR scheduling, where the maximum learning rate was set to 0.0001 multiplied by the batch size. LightStereo underwent training for 90 epochs. Only random crop ($320\times 736$) is employed. For the KITTI dataset, we fine-tuned the pre-trained models on the SceneFlow dataset [25] for 500 epochs using a mixed training set comprising KITTI 2012 [26] and KITTI2015 [27] training datasets. Batch size is set to 2. OneCycleLR scheduling is used with a max learning rate of 0.0002. For the generalization experiments, we employ data augmentation techniques including color jitter, random erase, random scale, and random crop. In the ablation study, these LightStereo models are trained over 50 epochs.
Figure 5: Qualitative results on SceneFlow [25].
TABLE II: Ablation study on SceneFlow [25] – Conv. Block Selection. We underline our final choice.
Conv. Type Kernel Block EPE (px) Flops(G) Param(M) Time(ms)
Regular $3\times 3$ (4 8 16) 0.7652 36.27 8.04 16.70
Regular $5\times 5$ (4 8 16) 0.7979 71.16 18.62 17.32
Regular $7\times 7$ (4 8 16) 0.8190 123.51 34.49 19.68
Regular $11\times 1$ 11 (4 8 16) 0.8672 280.53 82.10 33.60
V1 Block [20] DW $3\times 3$ (30 60 120) 0.7801 34.86 7.57 54.21
V2 Block [21] DW $3\times 3$ (4 8 16) 0.7144 35.82 7.54 22.93
ViT Block [40] - (3 6 9) 0.7149 34.48 6.53 51.14
@@ IV-B Implementation Details
TABLE III: Ablation study on SceneFlow [25] – Backbone selection. Comparison among efficient feature extractors.
@@
MobilenetV2 [21] CNN 0.7144 35.82 7.54 22.93
MobilenetV3 [48] CNN 0.7292 35.72 9.16 25.02
StarNet [49] CNN 0.7247 40.21 8.98 26.63
EfficientnetV2 [42] CNN 0.6130 103.14 28.87 46.83
RepVIT [50] Transformer 0.6823 50.45 9.56 28.65
Figure 6: Qualitative results on SceneFlow [25]. Comparison between LightStereo blocks.
TABLE IV: Ablation study on SceneFlow [25] – Block structure design. The last row refers to LightStereo-L.
Block Exp. factor SE [51] MSCA EPE (px) Flops(G) Param(M) Time(ms)
@@
TABLE V: Results on KITTI online benchmarks. Methods are classified based on runtime (higher or lower than 100ms). We highlight best and second best results. means time measured on our RTX 3090 GPU.
Method KITTI 2012[26] KITTI 2015[27] Time
3-all EPE-all D1-bg D1-fg D1-all (ms)
GANet[4] 1.60 0.91 1.48 3.46 1.81 1800
LaC+GANet[52] 1.42 0.80 1.44 2.83 1.67 1800
CFNet[53] 1.58 0.92 1.54 3.56 1.88 180
SegStereo[54] 2.03 1.25 1.88 4.07 2.25 600
SSPCVNet [55] 1.90 1.08 1.75 3.89 2.11 900
EdgeStereo-V2[56] 1.83 1.07 1.84 3.30 2.08 320
LEAStereo[7] 1.45 0.83 1.40 2.91 1.65 300
CREStereo[10] 1.46 0.90 1.45 2.86 1.69 410
ACVNet [12] 1.47 0.86 1.37 3.07 1.65 200
DispNetC[25] 4.11 2.77 4.32 4.41 4.34 60
StereoNet [43] - - 4.30 7.45 4.83 22
DeepPrunerFast[34] - - 2.32 3.91 2.59 50
AANet[18] 2.42 1.46 1.99 5.39 2.55 62
DecNet[57] - - 2.07 3.87 2.37 50
HITNet[45] 1.41 1.14 1.74 3.20 1.98 54
CoEx[39] 1.55 1.15 1.79 3.82 2.13 33
Fast-ACVNet[44] 1.68 1.23 1.82 3.93 2.17 39
Fast-ACVNet+[44] 1.45 1.06 1.70 3.53 2.01 45
LightStereo-S (Ours) 1.88 1.30 2.00 3.80 2.30 17
LightStereo-M (Ours) 1.56 1.10 1.81 3.22 2.04 23
LightStereo-L (Ours) 1.55 1.10 1.78 2.64 1.93 34
LightStereo-H (Ours) 1.34 0.96 1.60 2.92 1.82 49
TABLE VI: Runtime breakdown.
@@
LightStereo-H 27.17 1.98 23.64 1.49 54.28
TABLE VII: Generalization performance on multiple datasets. All the models are trained on the SceneFlow only.
Method KITTI12 D1(%) KITTI15 D1(%) Middle 2px(%)
DPF [34] 16.8 15.9 30.83
BGNet [36] 24.8 20.1 37.00
CoEx [39] 13.5 11.6 25.51
FastACV [44] 12.4 10.6 20.13
IINet [46] 11.6 8.5 19.57
LightStereo-S(Ours) 11.6 9.0 19.63
LightStereo-M(Ours) 7.0 6.6 17.69
LightStereo-L(Ours) 6.4 6.4 17.51
LightStereo-H(Ours) 7.2 7.3 14.27
IV-C Comparisons with State-of-the-art on SceneFlow
Table I compares LightStereo with several state-of-the-art stereo matching approaches on SceneFlow [25]. LightStereo-S runs in 17ms only, being substantially faster than other methods. Regarding model complexity, LightStereo-S strikes a favorable balance with only 22.71Gflops, comparable to StereoNet [43] (85.93Gflops). This reflects our commitment to maintaining a lightweight model while ensuring competitive performance. In terms of accuracy, LightStereo-H achieves an EPE of 0.51, resulting in more accuracy than Fast-ACVNet+ [44] (0.59) and IINet [46](0.54), yet remaining competitive in terms of complexity. Overall, our LightStereo framework, particularly the LightStereo-S configuration, presents a compelling solution for real-time stereo matching, offering a favorable trade-off between computational efficiency and depth estimation accuracy. Figure 5 presents a qualitative comparison of the results by the four proposed models.
IV-D Ablation Study
Conv. Block Selection. In Table II, we highlight the critical importance of disparity dimensions in the cost aggregation process of stereo matching, rather than spatial expansions in the height and width dimensions. On top, we can notice how regular convolutions with larger kernels yield both higher EPE and complexity. This finding indicates that merely increasing the spatial extent of convolutions does not effectively enhance the accuracy of stereo-matching; on the contrary, putting the focus on disparity dimensions is crucial. The V1 block with a $3\times 3$ kernel showed an EPE of 0.7801, indicating moderate performance with reduced FLOPs (34.86G), but it had a higher inference time (54.21ms). In contrast, the V2 block provided the best balance, achieving the lowest EPE of 0.7144, and FLOPs of 35.82G, with an inference time of 22.93ms. This superior performance is attributed to the structure of the V2 block, which incorporates expansion convolutions in the disparity dimension. The ViT block showed a similar EPE to the V2 block but had a lower parameter count (6.53M) and higher inference time (51.14ms). Ultimately, the V2 block was chosen for its optimal balance between accuracy and computational efficiency. Figure 6 shows a visual comparison with different conv. type. It can be observed that for the occluded area in the lower left corner of the left image, the cost aggregation based on the V2 block achieves better results than the other 3 conv. type.
Backbone Selection. Table III explores the use of classic, lightweight models as backbones. MobilenetV2 [21] demonstrates a balanced performance with an EPE of 0.7144, FLOPs of 35.82G, 7.54M parameters, and an inference time of 22.93ms. Although MobilenetV3 [48] had slightly lower FLOPs at 35.72G, it showed a higher EPE of 0.7292 and required more parameters (9.16M) with a longer inference time (25.02ms). StarNet [49] exhibited an EPE of 0.7247 with higher FLOPs (40.21G) and inference time (26.63ms). EfficientnetV2 [42] achieved the lowest EPE of 0.6130, but at the cost of significantly higher computational resources (103.14G FLOPs) and parameters (28.87M), with an inference time of 46.83ms. RepVIT [50], a transformer-based model, showed an EPE of 0.6876 but required 50.45GFLOPs and 9.56M parameters with an inference time of 28.65ms. Therefore, MobilenetV2 is chosen for its overall efficiency and balance between accuracy and computational cost.
Block Structure Analysis. In Table IV, we explore the impact of different block structures while maintaining a constant expansion factor of 4. Configurations ’e’ to ’h’ explore blocks (1, 2, 4), (2, 4, 8), (4, 8, 16), and (8, 16, 32) respectively. The results show that larger blocks generally lead to better performance. For instance, configuration ’e’ with the smallest block (1, 2, 4) has the highest EPE of 0.8317, while configuration ’h’ with the largest block (8, 16, 32) improves EPE, but detailed metrics are not provided. There is a trade-off between accuracy and computational cost, as larger blocks increase FLOPs and parameters.
Expansion Factor Analysis. Still in Table IV, we analyze the effect of varying expansion factors while keeping the block structure constant. As the expansion factor increases, a consistent decrease in EPE is observed, indicating improved accuracy. The analysis further proves that the critical importance of disparity dimensions cannot be overstated in the cost aggregation process of stereo matching. The disparity dimension is pivotal in effectively aggregating cost volumes, leading to more accurate and reliable depth estimations. However, this improvement in accuracy comes at the cost of significantly increased computational complexity and parameters, with flops increasing from 26.23G to 93.34G and the inference time increases from 22.44ms to 36.67ms.
MSCA Module Analysis. As illustrated in Table IV, the baseline configuration (i) with blocks (4, 8, 16) achieved an EPE of 0.7144. Incorporating MSCA (configuration k) reduced the EPE further to 0.6809 with minimal changes in FLOPs and parameters. This suggests that MSCA provides the most significant improvement in accuracy with minimal impact on computational efficiency. We also explored the use of the SE module [51]. Configurations ’i’ to ’j’ show that although the SE module improves accuracy, reducing EPE from 0.7144 to 0.7036. However, this improvement comes with an increase in parameters from 7.54M to 12.76M and a slight increase in FLOPs and inference time.